OpenStudy (toxicsugar22):

A doctor who is treated a heart patient wants to presicribe medication to lowere the patients blood pressure. The body's reaction to this medication is a function of the dose admininistered. If the patient takes x milliters of the medication, his blood pressure should decrease by r=f(x) points where f(x)=3x^2-1/3x^3 what values of x make sence for this function

OpenStudy (toxicsugar22):

can u please help

OpenStudy (anonymous):

Hmm.. I would say find the min value, and would that give you the x value? that would make the most sense?

OpenStudy (toxicsugar22):

i got 9 and 0

OpenStudy (anonymous):

Okay, Im not sure I understand the question, sorry o.O Because at x=0 f(x) =0 as well. I believe. I don't think that helps in anyway. Sorry

OpenStudy (amistre64):

y = 3x^2-1 ------- this? 3x^3

OpenStudy (anonymous):

Ohhh. I didn't realize that was division. haha

OpenStudy (anonymous):

sorry, but i dont know enough to help you out

OpenStudy (toxicsugar22):

amitre is that the answer

OpenStudy (amistre64):

no thats a question about the problem :) is that the question at hand?

OpenStudy (anonymous):

Any value greater than 0 would make sense. right..

OpenStudy (toxicsugar22):

no its what values of x make sence for this function

OpenStudy (amistre64):

i think we want the maximum for f(x) so that his blood pressure reduces the most right?

OpenStudy (amistre64):

well, any values of x will work except for x=0

OpenStudy (anonymous):

i think we are looking for domain ?

OpenStudy (amistre64):

y = 3x^2-1 ------- is this the function we are working with? 3x^3

OpenStudy (toxicsugar22):

yes

OpenStudy (amistre64):

then lets see where this function is increaseing and decreasing for x

OpenStudy (anonymous):

Domain should be any value greater than 0. But I don't think its that simple.

OpenStudy (anonymous):

no amistre has got it

OpenStudy (anonymous):

nned function to be decreasing, or blood pressure is not decreasing

OpenStudy (amistre64):

the 1st derivative is: 3x^3(2x) - 9x^2(3x^2-1) -21x^4 +9x^2 ------------- 9x^6

OpenStudy (amistre64):

when the derivative is zero we get the highes and lowes of it;

OpenStudy (amistre64):

-3x^2 (7x^2 -3) = 0 when x = 0 and... sqrt(12)/14 = sqrt(3)/7 the only value here that makes sense is sqrt(3/7)

OpenStudy (amistre64):

now is it a high or a low :) or an inflection point is there an answer key for this to determine if we are doing it right :)

OpenStudy (toxicsugar22):

no ther isnt

OpenStudy (amistre64):

3x^2-1 ------- 3x^3 3 sqrt.3/7^2-1 -------------- 3 sqrt.3/7^3 -40/49 -------------- thats a negative number so id say its not going to be that point :) 9/49 (sqrt.3/7)

OpenStudy (amistre64):

lets see what the roots are if any to this equation when 3x^2-1 =0 3x^2 = 1 x^2 = 1/3 x = sqrt(1/3) = sqrt(3)/3 is a zero

OpenStudy (toxicsugar22):

so that is the answer

OpenStudy (amistre64):

0.577 is a root

OpenStudy (amistre64):

dunno; still running tests in my head :)

OpenStudy (toxicsugar22):

ok its okay

OpenStudy (toxicsugar22):

can u help me with one more that is not realted to this

OpenStudy (toxicsugar22):

pomblem

OpenStudy (toxicsugar22):

please

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%283x^2-1%29%2F%283x^3%29 is what we get for a graph

OpenStudy (amistre64):

its only effective after x=sqrt(3)/3 and peaks at x=1

OpenStudy (toxicsugar22):

ok but can u help with anot

OpenStudy (amistre64):

i can try :)

OpenStudy (toxicsugar22):

ok it poynomials are u good at that

OpenStudy (amistre64):

i can be at times lol

OpenStudy (toxicsugar22):

Find the first three terms in asending powers. (Do not compute the entire product) (1-2x^2-x^4)(4+x^2-2x^4)

OpenStudy (amistre64):

4 -7x^2 -8x^4

OpenStudy (toxicsugar22):

thats the answer